soccer team raised 400.00 for equipment. will cost 41.50 for each of the 15 players How much more will each player have to pay?
Answer by vleith(2825)

Total = 41.50*15
Raised = 400
Total = raised + "whats left"
(41.50*15) - 400 = whatsleft
622.50 - 400 = whatsleft
222.50 = whatsleft
222.50/15 = whatEachPlayerMustPay
14.84

Or looking at another way
Each player costs 41.50
Each player's portion of the 400 = 400/15 = 26.66
41.50-26.66 = 14.84